OpenStudy (jacobbirddog):

The sum of three consecutive integers is –27. What are the numbers? A. {–11, –13, –15} B. {–5, –9, –13} C. {–10, –9, –8} D. {–7, –9, –11}

OpenStudy (anonymous):

Wanna play Minecraft with me?

OpenStudy (anonymous):

Faris doesn't give you answers. Lol.

OpenStudy (anonymous):

Lets consider n as the main number that u will add 1 and 2 to get next two values

OpenStudy (anonymous):

n+(n+1)+(n+2)=-27

OpenStudy (anonymous):

so it will be n+n+1+n+2=-27

OpenStudy (anonymous):

did u understand up to this point?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

It will be 3n+3=-27 Now solve n